# Active Inference AI Systems for Scientific Discovery

### v3

@@ Abstract
The rapid evolution of artificial intelligence has led to expectations of transformative impact on science, yet current systems remain fundamentally limited in enabling genuine scientific discovery. This perspective contends that progress turns on closing three mutually reinforcing gaps in abstraction, reasoning and empirical grounding. Central to addressing these gaps is recognizing complementary cognitive modes: thinking as slow, iterative hypothesis generation—exploring counterfactual spaces where physical laws can be temporarily violated to discover new patterns—and reasoning as fast, deterministic validation, traversing established knowledge graphs to test consistency with known principles. Abstractions in this loop should be manipulable models that enable counterfactual prediction, causal attribution, and refinement. Design principles—rather than a monolithic recipe— are proposed for systems that reason in imaginary spaces and learn from the world: causal, multimodal models for internal simulation; persistent, uncertainty-aware scientific memory that distinguishes hypotheses from established claims; formal verification pathways coupled to computations and experiments. It is also argued that the inherent ambiguity in feedback from simulations and experiments, and underlying uncertainties make human judgment indispensable, not as a temporary scaffold but as a permanent architectural component. Evaluations must assess the system’s ability to identify novel phenomena, propose falsifiable hypotheses, and efficiently guide experimental programs toward genuine discoveries.
@@
Over the past decade, the evolution of AI foundation model research has followed a clear sequence of discrete jumps in capability. The advent of the Transformer[62] marked a phase dominated by architectural innovations, which was rapidly succeeded by scaling demonstrations such as GPT-2[55]. The maturation of large-language-model pre-training then gave way to the usability turn: chat-oriented models fine-tuned for alignment and safety that enabled direct human interaction[47]. The current frontier is characterised by reasoning-emulation systems that incorporate tool use, scratch-pad planning, or program-synthesis objectives[45]. A fifth, still-incipient phase points toward autonomous agents that can decompose tasks, invoke external software or laboratories, and learn from the resulting feedback. Scientific applications of AI have echoed each of these transitions at a compressed cadence. As examples, SchNet translated architectural advances to quantum chemistry[59]; AlphaFold leveraged domain knowledge infused scaling to solve protein-fold prediction[30]; ChemBERTa [14] and FourCastNet [48] adapted language and vision innovations to molecular and climate domains; and AlphaGeometry applied reasoning-centric objectives to symbolic mathematics[61]. Collectively, recent works [24, 7, 9] chart a shift from single, specialized pre-trained model to workflow orchestration, suggesting that future breakthroughs may hinge on integrating heterogeneous, domain-aware agents capable of planning experiments, steering simulations, and iteratively refining hypotheses across scales.
@@
This highlights a deeper challenge for scientific discovery, which must reason across stacked layers of abstraction: the emergence of unexpected phenomena at higher scales, just as local atmospheric equations do not directly predict large-scale El Niño patterns. To address this challenge, it may be required to deliberately architect systems with built-in mechanisms for hierarchical inference, equipping them with specialized components that can navigate between reductionist details and emergent phenomena. A compelling counter-argument posits that such abstract reasoning is not a feature to be explicitly engineered, but an emergent property that will arise from sufficient scale and diverse data. Proponents of this view might point to tools such as AlphaGeometry [61], where complex, formal reasoning appears to emerge from a foundation model trained on vast synthetic data. However, we contend that while scaling can master any pattern present in the training distribution—even highly complex ones—it is fundamentally limited to learning correlational structures. Scientific discovery, in contrast, hinges on understanding interventional and counterfactual logic: what happens when the system is deliberately perturbed? This knowledge cannot be passively observed in static data; it must be actively acquired through interaction with the world or a reliable causal model thereof. The ’reality gap’ thus remains a significant barrier that pure scaling may not cross.
@@
It is also pertinent to examine the nature of present-day scientific discovery before speculating the role of AI. Modern science has moved beyond the romanticized vision of solitary geniuses grappling with nature’s mysteries. It may be difficult to generalize or even define the nature of discovery, but it is safe to assume that many of today’s discoveries emerge from vast collaborations parsing petabytes of data from instruments such as the Large Hadron Collider or from distributed sensor networks or large-scale computations and most importantly, refining hypothesis in concurrence with experiments and simulations. In fields such as high-energy physics, the bottleneck has shifted toward complexity management, whereas in data-constrained arenas such as fusion-plasma diagnostics, insight scarcity remains dominant; any general framework must therefore account for both regimes. Even if one possesses the raw data to answer profound questions, we often lack the cognitive architecture to navigate the combinatorial explosion of hypotheses, interactions, and emergent phenomena. This creates an opportunity for AI systems—to excel precisely where human cognition fails, in maintaining consistency across very high-dimensional parameter spaces, identifying and reasoning about subtle patterns in noisy data. At this juncture, it has to be emphasized that generating novel hypotheses might be the easy part [25]: the challenge is in rapidly assessing the impact of a hypothesis or action in an imaginary space. Thus AI systems have to be equipped with rich world models that can rapidly explore vast hypothesis spaces, and integrated with efficient computations and experiments to provide valuable feedback.
@@
Against this backdrop, this perspective piece is organized around three interlocking hurdles (i) the abstraction gap, which separates low-level statistical regularities from the mechanistic concepts on which scientists actually reason; (ii) the reasoning gap, which limits today’s models to correlation-driven pattern completion rather than causal, counterfactual inference; and (iii) the reality gap, which isolates computation from the empirical feedback loops that ultimately arbitrate truth. The central thesis is that scientific discovery demands a reimagining of AI architectures. Future AI systems must integrate active inference principles, maintaining persistent scientific memories while engaging in closed-loop interaction with both simulated and physical worlds. The following sections examine each gap in detail before presenting an integrated architecture that addresses these challenges holistically.
1   Fundamental Gaps in AI Models
The path from current AI capabilities to genuine scientific discovery is obstructed by interconnected barriers that reflect deep architectural limitations rather than scaling or engineering challenges. These gaps are not independent failures but symptoms of a unified problem: current AI systems lack the cognitive architecture necessary for scientific thinking. Understanding these gaps requires recognizing that they form a mutually reinforcing system of constraints: without rich abstractions there is little substrate for reasoning, and without tight coupling to reality even the most elegant abstractions may drift toward irrelevance.
1.1   The Abstraction Gap
@@
While early models largely manipulated tokens and pixels, recent advances in concept-bottleneck networks[35], symmetry-equivariant graph models[60], and neuro-symbolic hybrids[40] show preliminary evidence that contemporary AI can already represent and reason over higher-order scientific concepts and principles. Yet a physicist reasons in conservation laws and symmetry breaking, whereas language models still operate on surface statistics. Closing this abstraction gap requires addressing several intertwined weaknesses.
Inset 1: Thinking and reasoning Understanding how AI systems can enable scientific discovery requires examining the cognitive processes that these systems attempt to emulate. In this context, a critical distinction emerges between thinking and reasoning: Thinking can be operationalized as an iterative, exploratory process—searching for partial solutions in the form of patterns without guaranteed convergence. It is the slow, generative phase where new connections form and novel patterns emerge from a number of possibilities. Reasoning, by contrast, represents the fast, deterministic traversal of established knowledge structures—building the most expressive path through a graph of already-discovered patterns. This dichotomy [29] may explain why current AI systems excel at certain tasks while failing at others. Large language models can reason impressively when the requisite patterns exist in their training data—they rapidly traverse their learned knowledge graphs to construct seemingly intelligent responses. Yet they struggle with genuine thinking: the patient, iterative discovery of patterns that do not yet exist in their representational space. Scientific discovery demands both capabilities in careful balance. Thinking generates hypotheses by discovering new patterns through mental simulation and exploration; reasoning then rapidly tests these patterns against existing knowledge and empirical constraints. The purpose of thinking therefore is not to solve problems directly but to expand the pattern vocabulary available for subsequent reasoning. Each thinking cycle potentially adds new nodes and edges to the knowledge graph, creating shortcuts and abstractions that make previously intractable reasoning paths suddenly accessible. This is perhaps why breakthrough discoveries often seem obvious in retrospect—the thinking phase has restructured the problem space so thoroughly that the reasoning path becomes trivial. It is also possible to formalise the intuitive split between slower counterfactual thinking and faster deductive reasoning by mapping them onto the well-studied System-1/System-2 dichotomy in cognitive science [31, 21]. Recent neuro-symbolic RL agents already hint at this synergy: the survey of Acharya et al.[1] chronicles agents that fuse neural perception (System 1) with first-order symbolic planners (System 2) while Mao et al.[40] demonstrate compositional question-answering by training a neural concept learner that hands off logic programs to a symbolic executor.
@@
Modern transformer variants assemble chain-of-thought proofs[65] by replaying patterns observed during pre-training; they do not build explicit causal graphs or exploit formal logic engines except in narrow plug-in pipelines. As a result they fail at problems that demand deep compositionality. Several other shortcomings have also been pointed out [42].
@@
The gap between correlation and causation represents perhaps the most fundamental challenge in automated scientific discovery. While current models excel at finding statistical regularities, scientific understanding requires the ability to reason about interventions—to ask not just “what correlates with what?" but “what happens when we change this?". Pearl’s causal hierarchy [49] distinguishes three levels of cognitive ability: association (seeing), intervention (doing), and counterfactuals (imagining). Current AI systems operate primarily at the associative level, occasionally reaching intervention through experimental design. True scientific reasoning requires all three, particularly the counterfactual ability to imagine alternative scenarios that violate observed correlations. This connects directly to ethologist Konrad Lorenz’s insight –first tied to learning systems by Scholkopf [58]– that thinking is fundamentally about acting in imaginary spaces where we can violate the constraints of observed data. This mental experimentation—impossible in physical reality but accessible in the imagination—forms the basis of scientific law formation, as explained in Inset 1.
Critically, this mental experimentation must integrate diverse modalities of scientific data and representation. Scientific phenomena manifest as continuous vector/tensor fields over space-time (e.g., velocity-pressure fields, concentration gradients, electromagnetic fields) interleaved with discrete events (chemical reactions, phase transitions) and symbolic constructs (reaction mechanisms, theoretical frameworks). Effective AI systems for discovery must therefore maintain multi-modal embeddings that seamlessly translate between these representational levels—from raw sensor data to mathematical abstractions to causal hypotheses—enabling the system to reason simultaneously across observational patterns, physical mechanisms, and theoretical principles.

### v4

@@ Abstract
The rapid evolution of artificial intelligence has led to expectations of transformative scientific discovery, yet current systems remain fundamentally limited by their operational architectures, brittle reasoning mechanisms, and their separation from experimental reality. Building on earlier work–including foundational contributions in automated scientific discovery [58], closed-loop laboratory systems [54, 14, 92], and causal machine learning [75, 85]–this perspective contends that progress in AI-driven science now depends on closing three fundamental gaps—the abstraction gap, the reasoning gap, and the reality gap—rather than on model size/data/test time compute. While prior systems have made substantial progress on individual gaps, we argue that genuinely novel scientific discovery –defined here as the identification of previously unknown causal mechanisms, physical laws, or theoretical frameworks that generalize beyond training distributions– requires their integrated resolution. Scientific reasoning demands internal representations that support simulation of actions and response, causal structures that distinguish correlation from mechanism, and continuous calibration. Active inference AI systems for scientific discovery are defined as those that (i) maintain long-lived research memories grounded in causal self-supervised foundation models, (ii) employ symbolic or neuro-symbolic planners equipped with Bayesian guardrails, (iii) grow persistent knowledge graphs where thinking generates novel conceptual nodes, reasoning establishes causal edges, and real-world interaction prunes false connections while strengthening verified pathways, and (iv) refine their internal representations through closed-loop interaction with both high-fidelity simulators and automated laboratories—an operational loop where mental simulation guides action and empirical surprise reshapes understanding. In essence, this work outlines design principles for an architecture in which discovery arises from the interplay between internal models that enable counterfactual reasoning and external validation that grounds hypotheses in reality. It is also argued that the inherent ambiguity in feedback from simulations and experiments, and underlying uncertainties makes human judgment indispensable, not as a temporary scaffold but as a permanent architectural component.
1   Present day AI Systems and Scientific Discovery
@@
Over the past decade, the evolution of AI foundation model research has followed a clear sequence of discrete jumps in capability. The advent of the Transformer[96] marked a phase dominated by architectural innovations, which was rapidly succeeded by scaling demonstrations such as GPT-2[81]. The maturation of large-language-model pre-training then gave way to the “usability turn”: the shift from models optimized purely for benchmark performance to chat-oriented systems fine-tuned for alignment and safety that enabled direct human interaction[72]. The current frontier is characterised by reasoning-emulation systems that incorporate tool use, scratch-pad planning, or program-synthesis objectives[70]. A fifth, still-incipient phase points toward autonomous agents which are software systems capable of perceiving their environment, making decisions, and taking actions to achieve specified goals without continuous human supervision [103]that can decompose tasks, invoke external software or laboratories, and learn from the resulting feedback. Scientific applications of AI have echoed each of these transitions at a compressed cadence. As examples, SchNet translated architectural advances to quantum chemistry[87]; AlphaFold combined evolutionary multiple-sequence-alignment features with SE(3)-equivariant neural architectures [48] leveraging decades of accumulated protein structure data to achieve high accuracy, though still requiring petascale compute and vast unlabeled evolutionary sequences rather than dramatically fewer examples; ChemBERTa [23] and FourCastNet [74] adapted language and vision innovations to molecular and climate domains; and AlphaGeometry applied reasoning-centric objectives to symbolic mathematics[94]. Collectively, recent works [38, 9, 11] chart a shift from single, specialized pre-trained model to workflow orchestration, suggesting that future breakthroughs may hinge on integrating heterogeneous, domain-aware agents capable of planning experiments, steering simulations, and iteratively refining hypotheses across scales.
@@
This highlights a deeper challenge for scientific discovery, which must reason across stacked layers of abstraction: the emergence of unexpected phenomena at higher scales, just as local atmospheric equations do not directly predict large-scale El Niño patterns [20, 19]. This challenge of multi-scale abstraction has been recognized and discussed extensively in the automated scientific discovery literature [58]. To address this challenge, it may be required to deliberately architect systems with built-in mechanisms for hierarchical inference, equipping them with specialized components that can navigate between reductionist details and emergent phenomena. A compelling counter-argument posits that such abstract reasoning is not a feature to be explicitly engineered, but an emergent property that will arise from sufficient scale and diverse data. Proponents of this view might point to tools such as AlphaGeometry [94], where complex, formal reasoning appears to emerge from a foundation model trained on vast synthetic data. However, we contend that while scaling can master any pattern present in the training distribution—even highly complex ones—it is fundamentally limited to learning correlational structures. Scientific discovery, in contrast, hinges on understanding interventional and counterfactual logic [75, 43]: what happens when the system is deliberately perturbed? This knowledge cannot be passively observed in static data; it must be actively acquired through interaction with the world or a reliable causal model thereofthough we note that causal assumptions can sometimes be justified from observational data when appropriate structural constraints hold [18, 6]. The ‘reality gap’ thus remains a significant barrier that pure scaling may not cross.
@@
It is also pertinent to examine the nature of present-day scientific discovery before speculating the role of AI. Modern science has moved beyond the romanticized vision of solitary geniuses grappling with nature’s mysteries. It may be difficult to generalize or even define the nature of discovery, but it is safe to assume that many of today’s discoveries emerge from vast collaborations parsing petabytes of data from instruments such as the Large Hadron Collider or from distributed sensor networks or large-scale computations and most importantly, refining hypothesis in concurrence with experiments and simulations. In fields such as high-energy physics, the bottleneck has shifted toward complexity management, whereas in data-constrained arenas such as fusion-plasma diagnostics, insight scarcity remains dominant; any general framework must therefore account for both regimes. Even if one possesses the raw data to answer profound questions, we often lack the cognitive architecture to navigate the combinatorial explosion of hypotheses, interactions, and emergent phenomena. The connection between cognitive architecture and the capacity for rich abstraction is direct: effective reasoning over scientific concepts requires working memory structures that can hold, manipulate, and compose abstract representations [57], capabilities that current transformer architectures approximate but do not fully realize. This creates an opportunity for AI systems—to excel precisely where human cognition fails, in maintaining consistency across very high-dimensional parameter spaces, identifying and reasoning about subtle patterns in noisy data. At this juncture, it has to be emphasized that generating novel hypotheses might be the easy part [39]: the challenge is in rapidly assessing the impact of a hypothesis or action in an imaginary space. Thus AI systems have to be equipped with rich world models that can rapidly explore vast hypothesis spaces, and integrated with efficient computations and experiments to provide valuable feedback.
It is essential to distinguish between engineering discovery -the optimization of known systems toward well-defined objectives- and scientific discovery -the identification of novel causal mechanisms, physical laws, or theoretical frameworks. Engineering problems in materials science or drug design can often be addressed by closed-loop optimization systems operating over predefined search spaces [14, 92]. Scientific discovery, by contrast, requires the capacity for causal reasoning and theoretical insight that may restructure the problem space itself. Prior closed-loop systems have demonstrated genuine discoveries: Adam, an automated scientist for yeast metabolism, combined ontological reasoning with active learning to design and execute falsifiable experiments, leading to novel findings in genomics [54]; its successor Eve identified that triclosan is effective against malaria [100]. These systems embody the active inference principles we advocate: maintaining scientific memory while engaging in closed-loop interaction with physical laboratories. The contribution of this perspective is not to claim that such systems are impossible, but to argue that scaling beyond well-defined domains requires addressing the abstraction, reasoning, and reality gaps in an integrated manner, and to propose architectural principles for doing so.
Modern science also operates within a myriad of constraints that exist in economic, legal and social dimensions rather than physical laws. These constraints favor certain types of AI-driven discovery while effectively prohibiting others. Additionally, the structure of the modern scientific enterprise—with its emphasis on incremental, publishable units and citation metrics—may be fundamentally misaligned with the kind of patient, integrative thinking required for paradigm shifts. AI systems could theoretically ignore these social pressures, pursuing research programs too risky or long-term for humans. But this same freedom from social constraint raises the possibility of AI systems optimizing for discovery without broader goals accounted for. Perhaps more concerning, AI systems trained on existing scientific literature risk amplifying current biases and narrowing the space of explored ideas. In other words, AI systems might converge on well-studied paths [17, 65], reducing the rich variety of research directions to a handful of statistically probable avenues. Scientific progress demands not convergence but divergence—an explosion of hypotheses, methodologies, and frameworks that challenge orthodoxy. The challenge is in designing AI systems that expand rather than constrain the landscape of scientific imagination.
@@
Against this backdrop, the remainder of this perspective piece is organized around three interlocking hurdles that scientific discovery architectures must clear: (i) the abstraction gap, which separates low-level statistical regularities from the mechanistic concepts on which scientists actually reason; (ii) the reasoning gap, which limits today’s models to correlation-driven pattern completion rather than causal, counterfactual inference; and (iii) the reality gap, which isolates computation from the empirical feedback loops that ultimately arbitrate truth. Each gap both constrains and amplifies the others: without rich abstractions there is little substrate for reasoning, and without tight coupling to reality even the most elegant abstractions may drift toward irrelevance. While we treat these gaps in separate sections for expository clarity, they are deeply interconnected: the abstraction gap concerns what representations the system can manipulate, the reasoning gap concerns how those representations are transformed, and the reality gap concerns whether those transformations correspond to the external world. Closing one gap in isolation yields limited progress; for instance, rich abstractions without causal reasoning remain correlational, while causal reasoning without empirical grounding may drift into unfalsifiable speculation.
Philosophical Foundations The quest for scientific discovery via computation confronts a fundamental paradox. Gödel [37] proved that formal systems are incomplete, and cannot self-consistently prove all truths, while Wolfram [101] demonstrates that computational irreducibility pervades nature. Additionally, Penrose [76, 77] contends that human insight transcends algorithms. Yet scientific theories and computations are found to be highly effective in many cases. Insight can be gained from Wolfram’s recent comment [102]: The very presence of computational irreducibility necessarily implies that there must be pockets of computational reducibility, where at least certain things are regular and predictable. It is within these pockets of reducibility that science fundamentally lives. In most cases, these pockets cannot be deduced a priori—they require empirical discovery. This connects to Popper’s [78] falsificationism: we cannot prove we have found true reducibility, but we can discover boundaries through experiments that challenge assumptions. Empirical feedback escapes Gödel’s constraints while delineating where nature permits shortcuts. Kuhn’s analysis [56] adds temporal dynamics: Science alternates between normal science within established pockets and paradigm shifts that restructure understanding. AI systems must balance exploiting known regularities with flexibility to reconceptualize when evidence demands. This synthesis directly informs our architecture. Thinking explores for new pockets and tests boundaries; reasoning exploits discovered regularities. World models encode provisional maps of known pockets, subject to Popper’s falsification and Kuhn’s paradigm shifts. Human steering proves essential. Humans provide non-computational insight for recognizing genuine understanding, value judgments for directing exploration, and navigation through paradigm shifts where evaluation criteria themselves transform. Humans can shape the search process by encoding domain knowledge, identifying significant anomalies, and recognizing connections that form larger frameworks. When Faraday discovered electromagnetic induction, he did not deduce it from Maxwell’s equations (which did not yet exist)—he found it through experiment. Thus, productive collaborations can implement the complete scientific method: AI generates and tests hypotheses at scale; humans provide insight and judgment and empirical feedback provides critical steering.
2   The Abstraction Gap
@@
While early models largely manipulated tokens and pixels, recent advances in concept‐bottleneck networks[55], symmetry‐equivariant graph models[93], and neuro-symbolic hybrids[63] show preliminary evidence that contemporary AI can already represent and reason over higher‐order scientific concepts and principles. Yet a physicist reasons in conservation laws and symmetry breaking, whereas language models still operate on surface statistics. Closing this abstraction gap requires addressing several intertwined weaknesses.
@@
Modern transformer variants assemble chain-of-thought proofs[99] by replaying patterns observed during pre-training; they do not build explicit causal graphs or exploit formal logic engines except in narrow plug-in pipelines. As a result they fail at problems that demand deep compositionality. Several other shortcomings have also been pointed out [66]: (i) AI systems can create “illusions of understanding” where researchers mistake fluent outputs for genuine insight; (ii) they may narrow the diversity of research directions by channeling attention toward well-represented topics in training data; and (iii) they can erode researchers’ own reasoning skills through over-reliance on automated suggestions.
Thinking and reasoning A critical distinction emerges between thinking and reasoning: Thinking can be operationalized as an iterative, exploratory process—searching for partial solutions in the form of patterns without guaranteed convergence. It is the slow, generative phase where new connections form and novel patterns emerge from a number of possibilities. Reasoning, by contrast, represents the fast, deterministic traversal of established knowledge structures—building the most expressive path through a graph of already-discovered patterns. This dichotomy [47] may explain why current AI systems excel at certain tasks while failing at others. Large language models can reason impressively when the requisite patterns exist in their training data—they rapidly traverse their learned knowledge graphs to construct seemingly intelligent responses. Yet they struggle with genuine thinking: the patient, iterative discovery of patterns that do not yet exist in their representational space. Scientific discovery demands both capabilities in careful balance. Thinking generates hypotheses by discovering new patterns through mental simulation and exploration; reasoning then rapidly tests these patterns against existing knowledge and empirical constraints. The purpose of thinking therefore is not to solve problems directly but to expand the pattern vocabulary available for subsequent reasoning. Each thinking cycle potentially adds new nodes and edges to the knowledge graph, creating shortcuts and abstractions that make previously intractable reasoning paths suddenly accessible. This is perhaps why breakthrough discoveries often seem obvious in retrospect—the thinking phase has restructured the problem space so thoroughly that the reasoning path becomes trivial. We note that thiw terminology differs from Kahneman’s [49] System 1/System 2 framework, where System 2 denotes slow, deliberate reasoning. Our usage follows Johnson-Laird’s [47] mental models tradition, where “thinking” refers to the creative construction of new mental models (slow, exploratory), while “reasoning” refers to drawing inferences within established models (fast once patterns exist). The key insight is that scientific discovery requires both: the slow construction of novel conceptual frameworks and the rapid verification of their logical consequences. Both frameworks recognize the importance of dual cognitive modes; the difference lies in which dimension (effort vs. novelty) defines the slow/fast distinction. Recent neuro-symbolic RL agents hint at this synergy: the survey of Acharya et al.[1] chronicles agents that fuse neural perception with first-order symbolic planners while Mao et al.[63] demonstrate compositional question-answering by training a neural concept learner that hands off logic programs to a symbolic executor.
The gap between correlation and causation represents perhaps the most fundamental challenge in automated scientific discovery. While current models excel at finding statistical regularities, scientific understanding requires the ability to reason about interventions—to ask not just “what correlates with what?" but “what happens when we change this?"
@@
Pearl’s causal hierarchy [75] distinguishes three levels of cognitive ability: association (seeing), intervention (doing), and counterfactuals (imagining). Current AI systems operate primarily at the associative level, occasionally reaching intervention through experimental design. True scientific reasoning requires all three, particularly the counterfactual ability to imagine alternative scenarios that violate observed correlations. This claim finds support in the history of science: the link between smoking and cancer was established through epidemiological reasoning about counterfactuals without randomized controlled trials [27]. Similarly, Darwin’s theory of evolution was fundamentally a counterfactual framework: imagining what would happen to populations under different selection pressures. This connects directly to ethologist Konrad Lorenz’s insight [61] –first tied to learning systems by Scholkopf [85], building on Pearl’s foundational work [75]– that thinking is fundamentally about acting in imaginary spaces where we can violate the constraints of observed data. This mental experimentation—impossible in physical reality but accessible in the imagination—forms the basis of scientific law formation.
3   Unhobbling Intelligence
A certain level of consensus appears to be forming in the community that incremental scaling of present architectures may not deliver the qualitative leap that scientific discovery demands. Progress hinges on unhobbling—removing the design constraints that keep today’s models predictable, yet fundamentally limited—through concurrent advances in algorithms, speculation control, hardware co-design, and access models.
Algorithmic Gains
Future systems must balance the complementary modes of thinking and reasoning as first-class architectural principles. Thinking—or slow, iterative discovery of new patterns—demands (i) world-model agents that can explore counterfactual spaces through mental simulation [41]; (ii) curiosity-driven mechanisms that reward pattern novelty over immediate task performance [84, 73]; and (iii) mechanisms for preventing premature convergence, such as entropy regularization [42] or explicit exploration bonuses that maintain hypothesis diversity during search [7]. Reasoning—the fast, deterministic traversal of pattern graphs—demands (i) efficient knowledge graph architectures with learned traversal policies [29]; (ii) neuro-symbolic stacks that maintain both continuous representations and discrete logical structures[63]; and (iii) caching mechanisms that transform expensive thinking outcomes into rapid reasoning primitives. The interplay between these modes mirrors how scientists alternate between exploratory experimentation (thinking) and theoretical derivation (reasoning).
The notion that “thinking is acting in an imaginary space"—as Konrad Lorenz observed [61]—provides a foundational principle for understanding how world models enable scientific discovery. Just as biological organisms evolved the capacity to simulate actions internally before committing physical resources, AI systems with rich world models can explore vast hypothesis spaces through mental simulation. This capability enables a qualitatively different mode of inference compared to pure pattern matching: it supports counterfactual reasoning, experimental design optimization, and the anticipation of empirical surprises before they manifest in costly real-world experiments. World models can serve as the substrate for this imaginary action space, encoding not just correlations but causal structures that permit intervention and manipulation. The fidelity of these mental simulations—their alignment with physical reality—determines whether the system’s thoughts translate into valid discoveries.
Scientific progress thrives on disciplined risk: venturing beyond received wisdom while remaining falsifiable. Current alignment protocols deliberately dampen exploratory behaviour, biasing models toward safe completion of well-trodden trajectories. Controlled speculation frameworks—for example, curiosity-driven reinforcement learning [71] combined with Bayesian epistemic guards—could allow systems to seek novel hypotheses, flag them with calibrated uncertainty, and propose targeted experiments for arbitration. Mechanisms such as self-consistency voting [98], adversarial peer review, and tool-augmented chain-of-thought audits offer additional scaffolding to keep high-variance reasoning tethered to empirical reality.